Faced with rising fax costs, a firm issued a guideline that transmissions of 10 pages or more should be sent by 2-day mail instead. Exceptions are allowed, but they want the average to be 10 or below. The firm examined 35 randomly chosen fax transmissions during the next year, yielding a sample mean of 14.44 with a standard deviation of 4.45 pages. (a) At the .01 level of significance, is the true mean greater than 10? (b) Use Excel to find the right-tail p-value.
Answer by stanbon(75887)

Faced with rising fax costs, a firm issued a guideline that transmissions of 10 pages or more should be sent by 2-day mail instead. Exceptions are allowed, but they want the average to be 10 or below. The firm examined 35 randomly chosen fax transmissions during the next year, yielding a sample mean of 14.44 with a standard deviation of 4.45 pages.
(a) At the .01 level of significance, is the true mean greater than 10?
Ho: u <= 10
Ha: u > 10
-----------------
Test statistic:
t(14.44) = (14.44-10)/[4.45/sqrt(35)] = 5.9028
(b) Use Excel to find the right-tail p-value.
The p-value = tcdp(5.9028,100,df=34) = 0.0000005758...
==============
Cheers,
Stan H.